# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

### v2

@@ III-B1 Predicate Filter
Now, let $\mathcal{T}_{p}=\{(\tau,l)|l=1,(\tau,l)\in\mathcal{T}\}$ be the positive set of trajectories in $\mathcal{T}$ and $\mathcal{T}_{n}=\mathcal{T}\setminus\mathcal{T}_{p}$ be the negative set. For each predicate $\psi\in\Psi$, we then calculate $d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, and then calculate the cosine similarity $S$ between those distributions. We then remove predicates from the set $\Psi$ (and their negation from $\Psi_{\neg}$) which have a cosine similarity $S>S_{\mathrm{th}}$, where $S_{\mathrm{th}}$ is a user-defined threshold.
@@ III-B2 Loss Regularization
$R_{\mathrm{T}}(\mathbf{w}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}s_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (17)
Intuitively, this regularizer encourages the task and constraint clauses to contain different atomic predicates.
@@
where $\mathbf{w}^{\mathcal{O}}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$ defines the weights within one of our temporal operators and $\tilde{w}^{\mathcal{O}}_{ij},\tilde{w}^{\mathcal{O}}_{jk}\in\mathbf{w}^{\mathcal{O}}$. Intuitively, this regularizer encourages disjunctive clauses within a temporal clause to contain different predicates.
@@
~$\tilde{\mathcal{L}}(\tau,l,\mathbf{w})=\mathcal{L}(\tau,l,\mathbf{w})+$ $\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w})$
@@
where $\mathcal{L}$ is a classification loss and $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are regularizer hyperparameters. We use $\lambda_{R_{\mathrm{T}}}=0.01$ and $\lambda_{R_{\mathrm{D}}}=0.1$ for this work. We use the classification loss from [21],
@@
where $\zeta\in\mathbb{R}_{>0}$ is a hyperparameter ($\zeta=1$ for this work).
@@ III-B3 Weight Pruning
The weight pruning step further simplifies the currently inferred explanation by removing unnecessary weights. First, all weights with a value of zero are removed from the network. Then, among the remaining weights, the $N_{\mathrm{w}}\in\mathbb{Z}_{>0}$ smallest weights—where $N_{\mathrm{w}}$ is a user-specified hyperparameter—are removed to eliminate the least contributing parts of the specification. The remaining weights are then normalized. The pruning process terminates if there is only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$ or $\mathbf{w}^{\mathcal{G}}$ to ensure the task-constraint structure of the inferred explanation.
@@ III-D Explainability Evaluation Metrics
Beyond classification performance, TLNet also emphasizes the human-interpretability of inferred explanations. We quantify this emphasis by introducing three explainability metrics—conciseness, consistency, and strictness.
@@ III-D2 Consistency
$\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\sigma_{n}\in\Sigma_{n}}\sum_{\hat{\phi}_{n}\in\hat{\Phi}_{n}}\frac{\mathbf{1}_{\sigma_{n}}(\hat{\phi}_{n})}{K\lvert\Sigma_{n}\rvert}&\text{if $\Sigma_{n}\neq\emptyset$},\\ 0&\text{otherwise},\end{cases}$
where $\mathbf{1}_{\sigma_{n}}(\cdot)$ is the indicator function for $\sigma_{n}$. Consistency ranges from 0 to 1, with higher scores indicating greater agreement in clause structures across explanations, reflecting structural stability in learned explanations.
@@ III-D3 Strictness
Strictness ranges from 0 to 1, with higher scores implying stricter logical expressions by rewarding conjunctions and penalizing disjunctions.
@@ III-E Weight Distribution in wSTL
A key assumption in TLNet is that the weights can be distributed across Boolean and temporal clauses. We now prove that such a distribution is theoretically sound. The normalized weights of a wSTL formula in CNF can be syntactically distributed across Boolean operators as follows.
@@ IV Results
We perform a series of empirical experiments to compare our method to relevant baselines for robotic tasks. We consider three baselines for inference of wSTL explanations, each adapted to model our normalized wSTL structures (given by Equation 12): {outline} \1 NN-TLI [25]: this method implements a greedy pruning strategy that independently removes network weights that do not affect classification accuracy after optimization. We removed their weight-rounding process (to convert to STL) to consider the structure in Equation 12 and used the aggregation functions given in Figure 2, since our empirical results showed more training stability with those functions. \1 STONE (top-3) [21]: this method retains the 3 largest network weights after training. \1 STONE (top-5) [21, 24]: this method retains the 5 largest network weights after training. All baselines were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison. For each policy we explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy. Each dataset was split 80% and 20% for training and evaluation, respectively. Our code is available on GitHub after acceptance.
@@
Figure 4: Screenshots from our environments.
@@ IV-A1 Capture-the-flag (CtF)
CtF is a discrete grid-world environment with adversarial dynamics where two agents compete to capture each other’s flags, based on [19]. If both agents occupy adjacent cells in a given territory, the defending agent eliminates the other with 75% probability. We considered blue agent policies optimized for five scenarios. The first four considered different red agent heuristic policies (capture, fight, patrol, mixture) with the same reward function (capture: +1, kill: +0.5, being captured: $-1$, timestep penalty: $-0.01$). The fifth scenario (capture 0) used a capture red agent with the reward function modified to remove the kill reward (kill: 0). The red agent policies are defined as follows: capture pursues the flag; fight pursues the blue agent; patrol defends a boundary region; and mixture switches between capturing and defending. Red agents perform a random action 25% of the time. We defined eight atomic predicates based on the defeated (“df”) status $\{1,-1\}$ of the red agent and euclidean distances among the blue agent (“ba”), blue flag (“bf”), blue territory (“bt”), red agent (“ra”), red flag (“rf”), and obstacle (“ob”): $\psi_{\mathrm{ba,bt}}$, $\psi_{\mathrm{ba,ob}}$, $\psi_{\mathrm{ba,ra}}$, $\psi_{\mathrm{ba,rf}}$, $\psi_{\mathrm{ra,bf}}$, $\psi_{\mathrm{ra,bt}}$, $\psi_{\mathrm{ra,df}}$, $\psi_{\mathrm{ra,ob}}$.
@@ IV-A2 Fetch Push
We also consider a continuous control task based on the Fetch mobile manipulator [27]. The robot must push a blue block (“b”) to a red target (“t”) using its gripper (“g”), without disturbing a black obstacle (“o”). The state and action spaces are continuous ($s\in\mathbb{R}^{55},a\in\mathbb{R}^{4}$). Eight predicates were defined based on the distances between the gripper, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
@@ IV-A3 Robot Navigation
We included predicates irrelevant to task success in each environment (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) to evaluate whether the considered methods could eliminate them from their explanations.
@@ IV-B Policy Optimization
TLNet is a general framework to generate explanations of robot policies, though, we focused on policies optimized by deep reinforcement learning algorithms in this work. We used Stable-Baselines3 [29] for training all policies. For the CtF and navigation environments, PPO [30] was used; for the Fetch Push task, we used SAC [31] with hindsight experience replay (HER) [32]. Hyperparameters were taken from prior work [19, 33, 28]. For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explainability metrics. The time horizon ($H$) for each environment is 100 for CtF, 200 for Fetch Push, and 400 for Robot Navigation.
TABLE I: Comparison of the Most Frequently Inferred Explanations.
Scenarios TLNet (Ours) NN-TLI STONE (top-3) STONE (top-5)
CtF Capture
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.30\psi_{\mathrm{ba,rf}}\vee 0.70\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}_{I}^{\bar{w}}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33\neg\psi_{\mathrm{ra,bt}})\wedge$
$(0.09\neg\psi_{\mathrm{ba,bt}}\vee 0.07\neg\psi_{\mathrm{ra,bt}})]\wedge$
@@
$\mathcal{F}_{I}^{\bar{w}}[0.73\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.27\neg\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}_{I}^{\bar{w}}[0.83\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
CtF Fight
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.33\psi_{\mathrm{ba,rf}}\vee 0.67\neg\psi_{\mathrm{ba,ra}}]$
@@
$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]$
CtF Patrol
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.52\psi_{\mathrm{ba,rf}}\vee 0.48\neg\psi_{\mathrm{ra,bt}}]$
@@
$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ra,df}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ra,df}}]$
CtF Mixture
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.44\psi_{\mathrm{ba,rf}}\vee 0.56\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47\psi_{\mathrm{ba,rf}}$
$\wedge(0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.05\neg\psi_{\mathrm{ra,df}})\wedge(0.18\psi_{\mathrm{ba,rf}}\vee$
$0.03\psi_{\mathrm{ra,df}}\vee 0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ra,bt}})]$
$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}_{I}^{\bar{w}}[0.48\neg\psi_{\mathrm{ba,bt}}$
$\wedge 0.52\psi_{\mathrm{ba,rf}}]$
CtF Capture 0
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.31\psi_{\mathrm{ba,rf}}\vee 0.69\neg\psi_{\mathrm{ra,bf}}]$
@@
$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[(0.33\psi_{\mathrm{ba,rf}}$
$\vee 0.58\neg\psi_{\mathrm{ra,bf}})\wedge(0.05\psi_{\mathrm{ba,rf}}\vee 0.04\neg\psi_{\mathrm{ra,bt}})]$
$\mathcal{F}_{I}^{\bar{w}}[0.67\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.33\neg\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}_{I}^{\bar{w}}[0.83\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
Fetch Push
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$

### v3

@@ III-B1 Predicate Filter
Now, let $\mathcal{T}_{p}=\{(\tau,l)|l=1,(\tau,l)\in\mathcal{T}\}$ be the positive set of trajectories in $\mathcal{T}$ and $\mathcal{T}_{n}=\mathcal{T}\setminus\mathcal{T}_{p}$ be the negative set. For each predicate $\psi\in\Psi$, we compute $d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, and then calculate the cosine similarity $S$ between those distributions. We then remove predicates from the set $\Psi$ (and their negations from $\Psi_{\neg}$) which have a cosine similarity $S\geq S_{\mathrm{th}}$, where $S_{\mathrm{th}}$ is a user-defined threshold.
@@ III-B2 Loss Regularization
$R_{\mathrm{T}}(\mathbf{w}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (17)
Intuitively, this regularizer encourages the task and constraint clauses to contain different atomic predicates, and is independent of clause sparsity.
@@
where $\mathbf{w}^{\mathcal{O}}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$ defines the weights within one of our temporal operators and $\tilde{w}^{\mathcal{O}}_{ij},\tilde{w}^{\mathcal{O}}_{jk}\in\mathbf{w}^{\mathcal{O}}$. Intuitively, this regularizer encourages disjunctive clauses within a temporal clause to contain different predicates, and is again independent of clause sparsity.
@@
$\tilde{\mathcal{L}}(\tau,l,\mathbf{w})=\mathcal{L}(\tau,l,\mathbf{w})+$ $\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w})$
@@
where $\mathcal{L}$ is a classification loss and $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are regularizer hyperparameters. We use the classification loss from [27],
@@
where $\zeta\in\mathbb{R}_{>0}$ is a hyperparameter.
@@ III-B3 Weight Pruning
The weight pruning step further simplifies the currently inferred explanation by removing unnecessary weights. First, all weights with a value of zero are removed from the network. Then, among the remaining weights, the $N_{\mathrm{w}}\in\mathbb{Z}_{>0}$ smallest weights—where $N_{\mathrm{w}}$ is a user-specified hyperparameter—are removed to eliminate the least contributing parts of the specification. The remaining weights are then normalized. The pruning process is performed $N_{\mathrm{pr}}\in\mathbb{Z}_{>0}$ times and terminates if there is only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$ or $\mathbf{w}^{\mathcal{G}}$ to ensure the desired task-constraint structure. Empirically, increasing $N_{\mathrm{w}}$ leads to training instability, since pruning changes the computation graph structure of the neural network. Decreasing $N_{\mathrm{pr}}$ can lead to unnecessarily complex explanations.
@@ III-D Explainability Evaluation Metrics
Beyond classification performance, TLNet also emphasizes the human interpretability of inferred explanations. We quantify this emphasis by introducing three explainability metrics—conciseness, consistency, and strictness.
@@ III-D2 Consistency
$\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\sigma_{n}\in\Sigma_{n}}\sum_{\hat{\phi}_{n}\in\hat{\Phi}_{n}}\frac{\mathbb{I}(\sigma_{n}=\hat{\phi}_{n})}{K\lvert\Sigma_{n}\rvert}&\text{if $\Sigma_{n}\neq\emptyset$},\\ 0&\text{otherwise}.\end{cases}$
Consistency ranges from 0 to 1, with higher scores indicating greater agreement in clause structures across explanations, reflecting structural stability in learned explanations.
@@ III-D3 Strictness
Strictness ranges from 0 to 1, with higher scores implying stricter logical expressions by rewarding conjunctions and penalizing disjunctions. For example, given a temporal clause $\phi_{0}=\mathcal{F}[(0.5\psi_{0}\vee 0.25\psi_{1})\wedge 0.25\psi_{2}]$ and $P=6$, we get $C_{0}=1$ and $D_{0}=1$, with strictness of 0.167 for $\phi_{0}$.
@@ III-E Weight Distribution in wSTL
A key assumption in TLNet is that the weights can be distributed across Boolean and temporal clauses. We now prove that such a distribution is theoretically sound. The normalized weights of a wSTL formula in CNF can be syntactically distributed across Boolean operators as follows. [^1] Syntactic weight distributivity also holds for temporal operators by distributing operator weights to the normalized interval weights followed by distributions to the inner Boolean structures.
@@ IV Results
We perform a series of empirical experiments to compare our method to relevant baselines for robotic tasks. We consider three baselines for inference of wSTL explanations, each adapted to model our normalized wSTL structures (given by Equation 12):
  • NN-TLI [15]: this method implements a greedy pruning strategy that independently removes network weights that do not affect classification accuracy after optimization. We removed their weight-rounding process (to convert to STL) to consider the structure in Equation 12 and used the aggregation functions given in Equations 7, 8, 9 and 10, since our empirical results showed more training stability with those functions.
  • STONE (top-3, top-5) [27, 28]: this method retains the 3 or 5 largest network weights after training.
  • DT (STL) [6]: this method uses decision trees to mine STL (not wSTL) explanations.
The first two baselines, NN-TLI and STONE (top-3, top-5), were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison. For each policy we explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy. Each dataset was split 80% and 20% for training and evaluation, respectively. Hyperparameter values are summarized in Table I. Our code, including videos of target policies, is available online [^2] https://github.com/LIRA-illinois/tlnet.
@@
Figure 4: Initial states of our test environments with relevant objects shown.
TABLE I: Hyperparameter values and their ranges.
Environments $\sigma$ $\zeta$ $S_{\mathrm{th}}$ $\lambda_{R_{\mathrm{T}}}$ $\lambda_{R_{\mathrm{D}}}$ $N_{\mathrm{pr}}$ $N_{\mathrm{w}}$ $H$
CtF 0.5 1.0 0.99 0.01 0.1 34 1 100
Fetch Push 0.5 1.0 0.99 0.01 0.1 20 1 200
Robot Nav. 0.5 1.0 0.99 0.01 0.1 28 1 400
Range ($\in$) $\mathbb{R}_{>0}$ $\mathbb{R}_{>0}$ [-1, 1] $\mathbb{R}_{\geq 0}$ $\mathbb{R}_{\geq 0}$ $\mathbb{Z}_{>0}$ $\mathbb{Z}_{>0}$ $\mathbb{Z}_{>0}$
@@ IV-A1 Capture-the-flag (CtF)
CtF is a discrete grid-world environment with adversarial dynamics where two agents compete to capture each other’s flags, based on [29]. If both agents occupy adjacent cells in a given territory, the defending agent eliminates the other with 75% probability. We considered blue agent policies optimized for five scenarios. The first four considered different red agent heuristic policies (capture, fight, patrol, mix) with the same reward function (capture: +1, kill: +0.5, being captured: 1-1, timestep penalty: 0.01-0.01). The fifth scenario (capture 0) used a capture red agent with the reward function modified to remove the kill reward (kill: 0). The red agent policies are defined as follows: capture pursues the flag; fight pursues the blue agent; patrol defends a boundary region; and mix uses fight when the blue agent is in the red territory and capture otherwise. Red agents perform a random action 25% of the time. We defined eight atomic predicates based on the defeated (“df”) status $\{1,-1\}$ of the red agent and Euclidean distances among the blue agent (“ba”), blue flag (“bf”), blue territory (“bt”), red agent (“ra”), red flag (“rf”), and obstacle (“ob”): $\psi_{\mathrm{ba,bt}}$, $\psi_{\mathrm{ba,ob}}$, $\psi_{\mathrm{ba,ra}}$, $\psi_{\mathrm{ba,rf}}$, $\psi_{\mathrm{ra,bf}}$, $\psi_{\mathrm{ra,bt}}$, $\psi_{\mathrm{ra,df}}$, $\psi_{\mathrm{ra,ob}}$.
@@ IV-A2 Fetch Push
We also consider a continuous control task based on the Fetch mobile manipulator [9]. The robot must push a blue block (“b”) to a red target (“t”) using its grip (“g”), without disturbing a black obstacle (“o”). The state and action spaces are continuous ($s\in\mathbb{R}^{55},a\in\mathbb{R}^{4}$). Eight predicates were defined based on the distances between the grip, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
@@ IV-A3 Robot Navigation
We included predicates irrelevant to task success in each environment (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) to evaluate whether the considered methods could eliminate them from their explanations. All predicates are distance-based (e.g., $d_{\mathrm{ba,ra}}<0.5$ ) except for $\psi_{\mathrm{ra,df}}$ (binary). Average TLNet optimization times were 170.8 s (CtF), 118.3 s (Fetch Push), and 116.1 s (Robot Navigation). [^3] With an Intel Xeon Silver 4314 CPU and NVIDIA RTX A4000 GPU.
@@ IV-B Policy Optimization
TLNet is a general framework to generate explanations of robot policies, though, we focused on policies optimized by deep reinforcement learning algorithms in this work. We used Stable-Baselines3 [21] for training all policies. For the CtF and navigation environments, we used PPO [24]; for the Fetch Push task, we used SAC [11] with hindsight experience replay (HER) [2]. Hyperparameters were taken from prior work [29, 22, 13]. For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explainability metrics.
TABLE II: Comparison of the Most Frequently Inferred Explanations with Frequencies (%).
Scenario TLNet (Ours) NN-TLI STONE (top-3) STONE (top-5) DT
CtF
Capture
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.30\psi_{\mathrm{ba,rf}}\vee$
   $0.70\neg\psi_{\mathrm{ra,bf}}]$
60%
$0.5\mathcal{F}_{I}^{\bar{w}}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33$
$\neg\psi_{\mathrm{ra,bt}})\wedge(0.09\neg\psi_{\mathrm{ba,bt}}\vee 0.07\neg\psi_{\mathrm{ra,bt}})]\wedge$
@@
20%
$\mathcal{F}_{I}^{\bar{w}}[0.73\neg$
$\psi_{\mathrm{ra,bt}}\wedge 0.$
$27\neg\psi_{\mathrm{ba,bt}}]$
50%
$\mathcal{F}_{I}^{\bar{w}}[0.83\neg$
$\psi_{\mathrm{ra,bt}}\wedge 0$
$17\psi_{\mathrm{ba,rf}}]$
60%
$\mathcal{F}_{I}[\psi_{\mathrm{ba,rf}}$
$\vee\psi_{\mathrm{ra,bf}}\vee$
$\neg\psi_{\mathrm{ba,bt}}]$
100%
CtF
Fight
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.33\psi_{\mathrm{ba,rf}}\vee$
   $0.67\neg\psi_{\mathrm{ba,ra}}]$
70%
@@
70%
$\mathcal{F}_{I}^{\bar{w}}$
$1.0\psi_{\mathrm{ba,rf}}]$
100%
$\mathcal{F}_{I}^{\bar{w}}$
$1.0\psi_{\mathrm{ba,rf}}]$
100%
$\mathcal{F}_{I}[\psi_{\mathrm{ba,rf}}]$
$\wedge$
$\mathcal{G}_{I}[\psi_{\mathrm{ba,bt}}]$
100%
CtF
Patrol
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.52\psi_{\mathrm{ba,rf}}\vee$
   $0.48\neg\psi_{\mathrm{ra,bt}}]$
70%
@@
30%
$\mathcal{F}_{I}^{\bar{w}}$
$1.0\psi_{\mathrm{ra,df}}]$
100%
$\mathcal{F}_{I}^{\bar{w}}$
$1.0\psi_{\mathrm{ra,df}}]$
90%
$\mathcal{G}_{I}[\neg\psi_{\mathrm{ba,rf}}$
$\wedge\psi_{\mathrm{ba,bt}}]$
100%
CtF
Mix
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.44\psi_{\mathrm{ba,rf}}\vee$
   $0.56\neg\psi_{\mathrm{ra,bf}}]$
60%
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47$
$\psi_{\mathrm{ba,rf}}\wedge(0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.05\neg\psi_{\mathrm{ra,df}})\wedge(0.18\psi$
$_{\mathrm{ba,rf}}\vee 0.03\psi_{\mathrm{ra,df}}\vee 0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ra,bt}})]$
10%
$\mathcal{F}_{I}^{\bar{w}}$
$1.0\psi_{\mathrm{ba,bt}}]$
60%
$\mathcal{F}_{I}^{\bar{w}}[0.48\neg$
$\psi_{\mathrm{ba,bt}}\wedge 0$
$52\psi_{\mathrm{ba,rf}}]$
40%
$\mathcal{F}_{I}[\neg\psi_{\mathrm{ba,rf}}$
$\vee\neg\psi_{\mathrm{ba,ra}}]$
$\wedge\mathcal{G}_{I}[\psi_{\mathrm{ba,rf}}]$
100%
CtF
Capture
0
@@
$0.5\mathcal{G}_{I}^{\bar{w}}[0.31\psi_{\mathrm{ba,rf}}\vee$
   $0.69\neg\psi_{\mathrm{ra,bf}}]$
80%
@@
$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[(0.33\psi_{\mathrm{ba,}}$
$_{\mathrm{rf}}\vee 0.58\neg\psi_{\mathrm{ra,bf}})\wedge(0.05\psi_{\mathrm{ba,rf}}\vee 0.04\neg\psi_{\mathrm{ra,bt}})]$
20%
$\mathcal{F}_{I}^{\bar{w}}[0.67\neg$
$\psi_{\mathrm{ra,bt}}\wedge 0.$
$33\neg\psi_{\mathrm{ba,bt}}]$
60%
$\mathcal{F}_{I}^{\bar{w}}[0.83\neg$
$\psi_{\mathrm{ra,bt}}\wedge 0$
$17\psi_{\mathrm{ba,rf}}]$
40%
$\mathcal{F}_{I}[\psi_{\mathrm{ba,rf}}\vee$
$\neg\psi_{\mathrm{ba,ra}}\vee$
$\neg\psi_{\mathrm{ba,bt}}]$
100%
Fetch
Push
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]\wedge 0.5$
$\mathcal{G}_{I}^{\bar{w}}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
100%